Any help not good at these problems

J. Bond and E. Blofeld are driving speedboats in opposite directions after a failed attempt by the latter to blow up the world with a nuclear bomb. Mr. Bond’s boat is going 25 MPH faster than Mr. Blofeld’s. After 3 hours they are 500 miles apart. How fast was each of them going?
Eric

Answer by richwmiller(17219)

Have you ever heard of a good problem? They are called problems for a reason.
6x+75=500
6x=425
x=425/6=70.833 and x+25=95.833